OpenStudy (anonymous):

The distribution of images on the pages of a 350-page catalog is found to follow normal distribution with a mean of 32 images per page and a standard deviation of 8 images per page. Which conclusion does this information best support?

OpenStudy (anonymous):

A.Approximately 238 pages fall within the first standard deviation of the mean. B.Approximately 238 pages fall within the second standard deviation of the mean. C.Approximately 238 pages fall within the third standard deviation of the mean. D.Approximately 346 pages fall within the second standard deviation of the mean. E.Approximately 332 pages fall within the third standard deviation of the mean.

OpenStudy (amistre64):

we need to know what: x/350 is equal to

OpenStudy (anonymous):

32?

OpenStudy (amistre64):

each option has a number of pages value .... we need to calculate the percentage out of 350 pages total

OpenStudy (anonymous):

so do we plug in 238 first?

OpenStudy (amistre64):

sure, we want to know what 238/350 is, and also 246/350, and 332/350

OpenStudy (anonymous):

238/350=.68

OpenStudy (anonymous):

246/350=.70

OpenStudy (anonymous):

332/350= .94

OpenStudy (anonymous):

what do i do now?

OpenStudy (amistre64):

thats spose to be 346 we get: 68%, 99%, and 95% which follow the empirical rule

OpenStudy (anonymous):

346/350 = .98

OpenStudy (anonymous):